# Detector Response to Gravitational Wave Polarizations in Gravitational Quantum Field Theory

### v1

@@ Abstract
We present an analysis of gravitational wave polarization modes within Gravitational Quantum Field Theory (GQFT), a unified theoretical framework reconciling general relativity and quantum field theory. Our study focuses on five fundamental polarization states predicted in GQFT: two tensor (+,×+$\times$+ , ×), two vector (x, y), and one scalar (breathing) mode, focusing on their distinctive detection signatures in space-based interferometers like LISA and Taiji. Using first-order orbital dynamics in the Solar System Barycenter frame, we identify three novel observational features: (1) characteristic interference patterns between polarization modes, (2) distinctive null-point signatures enabling mode discrimination, and (3) sky-position-dependent optimal detection windows. Our approach provides complete sky coverage through polarization mapping while remaining fully compatible with existing mission designs, notably avoiding the need for challenging direct breathing-mode measurements. The results are presented through comprehensive sky maps, offering both theoretical insights into gravitational wave polarization and practical tools for future detector networks. This work establishes a new paradigm for testing fundamental gravity theories through their unique polarization fingerprints, with particular relevance for upcoming multi-messenger gravitational wave astronomy.
1 Introduction
Gravitational waves (GWs) have been extensively observed through various experiments, with polarization playing a crucial role in exploring extended theories of gravity. In the framework of metric-compatible theories, there are up to six possible polarization modes Eardley et al. (1973): two tensor modes (+ and $\times$), two vector modes (x and y), and two scalar modes (breathing and longitudinal). Different gravitational theories predict distinct polarization states. For instance, in general relativity, only the plus and cross modes exist, which have been detected by the LIGO, Virgo, and KAGRA collaborations Abbott et al. (2023). In Brans-Dicke gravity Brans & Dicke (1961); Brans (1962), an additional breathing mode arises due to the presence of a scalar field. Among various alternative theories, this paper focuses on Gravitational Quantum Field Theory (GQFT) Wu (2015, 2016, 2023, 2024).
GQFT was established to reconcile general relativity (GR) and quantum field theory (QFT). It is based on the fundamental principle that the laws of nature are determined by the intrinsic properties of matter’s basic constituents. This principle distinguishes between two types of symmetries: intrinsic symmetries, which are defined by the quantum numbers of quantum fields as elementary particles, and external symmetries, which describe their motion in flat Minkowski spacetimeWu (2015, 2016). According to this framework, the intrinsic spin symmetry SP(1,3) must be localized into a spin gauge symmetry following the gauge symmetry principle.
@@ 1 Introduction
The gravigauge field $\chi_{\mu}^{\;a}$ behaves as a Goldstone-boson- type entity, corresponding to a massless graviton. Similar to how gravitational interactions occur for Dirac fermions, the coupling between the gravigauge field $\chi_{\mu}^{\;a}(x)$ and the spin gauge field ${\mathcal{A}}_{\mu}^{ab}$ enables the construction of a spin- gauge-invariant action in GQFT. This structure cannot be replicated using the metric field alone. Importantly, the general linear group symmetry $\mathbb{R}$), which is fundamental to GR, appears as an implicit symmetry within GQFT. A key distinction from GR is that GQFT maintains a flat Minkowski spacetime as its underlying spacetime framework. Interactions between the gravigauge field and the spin gauge/ specified spinor fields introduce non-geometric effects, which result in violations of the equivalence principle observed in GR and give rise to novel physical phenomena Gao et al. (2024). Classically, GQFT recovers GR when the internal spin-gauge sector is deactivated and the new effects stemming from the equivalence principle violation are disregarded.
@@
Beyond its foundational implications, GQFT provides a theoretical framework for exploring a hyperunified field theory capable of unifying all fundamental interactions Wu (2017, 2018a, 2018b, 2021a, 2021b, 2022). Recently, a general theory of the standard model Wu (2025) has been developed within the GQFT framework. This theory integrates the two standard models of particle physics and cosmology, offering new insights into the mysteries of the universe’s dark sector. Specifically, it provides fresh perspectives on understanding the nature of dark matter, the dynamics of the early inflationary universe, and the role of dark energy in the current cosmic expansion.
@@
Numerous studies have been conducted within the framework of GQFT, addressing topics such as inflation Wang et al. (2023), dark matter Wang et al. (2022), and particle physics Huang et al. (2023). The Taiji mission, a space-based gravitational wave (GW) observatory, has been analyzed within the parametrized post-Einsteinian framework to evaluate its capability in detecting different polarization modes. Studies indicate that Taiji can measure both dipole and quadrupole GW emissions Collaboration (2021); Gong et al. (2021); Liu et al. (2020). Similarly, the LISA-Tianqin network has been investigated for its sensitivity to the polarization modes of the stochastic GW background Hu et al. (2024). In a general four-dimensional metric theory of gravity, the detection of additional polarization modes beyond the two predicted by GR would suggest the need for an extended theory of gravitation. The specific polarizations observed could help rule out certain theoretical models Nishizawa et al. (2009a). Both Taiji Hu & Wu (2017) and LISA Danzmann & team (1996) employ a triangular three-detector configuration with equal arm lengths. This paper focuses on the polarization modes in GQFT and their corresponding response characteristics in such a three-detector setup.
@@
The interaction of gravitational waves (GWs) with a detector arm is manifested as fluctuations in the arm length between two detectors in a triangular three-detector configuration. The intensity of the detector’s response depends on the angle between the GW propagation direction and the detector arm. This response is minimal when the GW propagation direction is parallel to the arm and reaches its maximum when the GW direction is perpendicular to the arm. Since gravitational wave sources are distributed across the sky, the angle between the GW propagation direction and the detector arm is not constant. This variation arises due to the orbital motion of space-based detectors, such as Taiji Hu & Wu (2017) and LISA Danzmann & team (1996), around the Sun. As these detectors continuously change their positions and orientations, the angle between the GW propagation direction and the detector arm fluctuates, leading to variations in the detector’s sensitivity to GW signals.
@@
Laser interferometers detect gravitational waves (GWs) by measuring the differential length changes between two detector arms. However, maintaining perfectly equal arm lengths over the entire orbital period of a space-based detector is impractical. This challenge necessitates the use of Time Delay Interferometry (TDI) to account for varying arm lengths. Recent advancements in atomic clock technology offer an alternative approach. By precisely measuring the optical path time difference between detectors, these clocks can infer arm-length variations without relying solely on TDI. This method has sparked significant interest in the development of high-precision optical clocks for space-based GW detectionKolkowitz et al. (2016); Tino et al. (2019); He & Zhang (2020); Wang et al. (2025).
@@
In this paper, we concentrate on investigating novel polarization modes predicted by GQFT that extend beyond GR. Our analysis specifically examines the response of individual detector arms to gravitational wave signals, and will not include considerations of dual-arm interferometric measurements to maintain analytical focus. The paper is organized as follows: it first introduces in §2.1 the polarization modes in metric theories of gravity, then derives in §2.2 the specific polarization modes within the GQFT framework. The analysis further presents in §2.3 the detector response characteristics in the Solar System Barycenter frame, along with a method to distinguish between GQFT and GR based on zero-point distribution patterns, and providing the optimal observational timing windows in §3.
2 Methods
2.1 Polarization in metric theory
@@
The action of GW on test masses can be understood through their tidal effects, which manifest as deviations from geodesic motion. This phenomenon is captured by the Jacobi field equation for geodesic deviation, $\dfrac{\mathop{}\!\mathrm{D}^{2}X^{\alpha}}{\mathop{}\!\mathrm{d}\tau^{2}}=T^{\gamma}\nabla_{\gamma}(T^{\delta}\nabla_{\delta}X^{\alpha})=-R^{\alpha}_{\beta\delta\gamma}X^{\delta}T^{\gamma}T^{\beta}.$ In metric theories of gravity, the polarization state is entirely determined by the electric components of the Riemann tensor Dadhich (2000). When examined in the gravitationally independent flat frame (the natural framework for GQFT), this description simplifies to the form, $\dfrac{\mathop{}\!\mathrm{d}^{2}X^{i}}{\mathop{}\!\mathrm{d}\tau^{2}}=-R^{i}_{0j0}X^{j}$, which describes weak, plane, null gravitational waves Eardley et al. (1973). The symmetry properties of the Riemann tensor reveal that $R^{i}_{0j0}$ possesses six independent components, indicating that gravitational waves can exhibit up to six distinct polarization modes.
@@
For a wave propagating along the z-direction, we define $R_{i0j0}(t)$ as the “driving force matrix” $S_{ij}(t)$, a 3-dimensional symmetric matrix. The polarization modes can be further classified using the Newman-Penrose formalism Newman & Penrose (1962). Working with tetrad indices $(l,n,m,\bar{m})$, we consider tensor contractions with the corresponding null vectors, $X_{efgh}:=X_{\mu\nu\rho\sigma}e^{\mu}f^{\nu}g^{\rho}h^{\sigma}$, where $(e,f,g,h)$ range over the null tetrad basis $(l,n,m,\bar{m})$. Taking a null tetrad frame as follows:
$l=\frac{1}{\sqrt{2}}(\partial_{t}+\partial_{z}),\quad n=\frac{1}{\sqrt{2}}(\partial_{t}-\partial_{z}),$
$m=\frac{1}{\sqrt{2}}(\partial_{x}+i\partial_{y}),\quad\bar{m}=\frac{1}{\sqrt{2}}(\partial_{x}-i\partial_{y}).$ (1)
@@
In the null frame, where the field depends solely on retarded time $(t-z)$, the Riemann tensor satisfies the condition, $R_{efgh,p}=0$. Here, $(e,f,g,h)$ span the null tetrad $(l,n,m,\bar{m})$, while $(p,q)$ range over $(l,m,\bar{m})$. Substituting these into the differential Bianchi identity yields, $R_{ef[pq,n]}={\dfrac{1}{3}R_{efpq,n}=0}$. This implies that, up to trivial constants, we must have: $R_{efpq}=R_{pqef}=0$. Consequently, the only non-vanishing components of the Riemann tensor are of the form $R_{pnqn}$.
@@
Under the null condition, the original 12 NP scalars reduce to just 4 independent quantities: $\Psi_{2},\Psi_{3},\Psi_{4},\Phi_{22}$. Two real scalars $\Psi_{2},\Phi_{22}$, and Two complex scalars $\Psi_{3},\Psi_{4}$. These remaining NP null scalars collectively contain 6 independent degrees of freedom, corresponding exactly to the degrees of freedom in the Riemann tensor for weak, plane, null gravitational waves Hyun et al. (2019):
@@
where $A$ ranges over (+,×$(+,\times,x,y,b,l),$ covering 6 polarization modes.
@@ 2.2 Polarization in GQFT
Dynamical analysis of GQFT identifies five fundamental degrees of freedom in the gravitational sector Gao et al. (2024). Although this implies the theoretical existence of five gravitational wave polarization modes, only three physical polarizations emerge as observationally significant when analyzing their tidal effects on test masses. The geodesic deviation equation, which governs observable gravitational wave effects, exclusively captures tidal forces. Consequently, the remaining two vector-mode polarizations remain observationally inaccessible within this framework, as they encode intrinsic spin properties and represent manifestations of equivalence principle violations rather than measurable tidal deformations.
@@
[$\ ~{}&\hat{h}_{+}&\hat{h}_{\times}&~{}\\ ~{}&\hat{h}_{\times}&-\hat{h}_{+}&~{}\\ ~{}&~{}&~{}&~{}\end{bmatrix},$ (6)
@@
2.3 Response on the detectors
As mentioned above, GW propagation can be expressed as a linear combination of different polarization tensors Nishizawa et al. (2009b),
@@ 2.3 Response on the detectors
Assuming the tensor mode is dominated by quadrupole radiation while other modes exhibit dipole-dominated radiation, the waveform incorporating different polarizations is expressed within the parametrized post-Einsteinian framework at 1PN order Epstein & Wagoner (1975); Wagoner & Will (1976); Will (1977); Chatziioannou et al. (2012). For simplicity, we neglect the orbital parameters of the wave source, considering only the waveform amplitudes. Since this work aims to develop a polarization-focused model for distinguishing between GQFT and GR, the analysis remains independent of specific waveform models, thereby maintaining generality.
A comparative analysis of GQFT wave equations reveals that the b-mode and tensor modes differ by a characteristic factor $(1+\gamma_{W})$, where $\gamma_{W}$ represents a fundamental GQFT parameter characterized by the mass ratio $\gamma_{W}=M_{\mathcal{A}}^{2}/\bar{M}_{\kappa}^{2}$ Wu (2015, 2016, 2023); Gao et al. (2024); Wu (2024, 2025). Here, $M_{\mathcal{A}}$ is the mass of spin gauge boson and $\bar{M}_{\kappa}$ denotes the fundamental mass at Planck scale. As previously established, the polarization amplitudes for plus, cross, and other modes are given by: $\dfrac{2\mathcal{M}}{d_{L}}(\mathcal{M}\omega)^{\frac{2}{3}}$, $\dfrac{4\mathcal{M}}{d_{L}}(\mathcal{M}\omega)^{\frac{2}{3}}$ and $\dfrac{\alpha_{A}\mathcal{M}}{d_{L}}(\mathcal{M}\omega)^{\frac{1}{3}}$ respectively Liu et al. (2020); O’Beirne et al. (2019). Here, $d_{L}$ is the luminosity distance, $\mathcal{M}=M\eta^{\frac{3}{5}}$ is the chirp mass with $\eta=\dfrac{m_{1}m_{2}}{M^{2}}$ and $M=m_{1}+m_{2}$ is the total mass.
Turning to detector trajectories, both the LISA and Taiji missions have Keplerian orbital characteristics, maintaining an arm length of 2.5$\sim$ 3 million kilometers. To first order in eccentricity, the spacecraft positions in SSB coordinates can be expressed as Rubbo et al. (2004),
@@
When examining the effect of GWs on the detector, consider a photon emitted from spacecraft $i$ at time $t_{i}$ and received at spacecraft $j$ at time $t_{j}$. The nominal arm length is $\ell_{ij}$, and its variation under low-frequency GWs (where the transfer function approaches unity) can be expressed as Rubbo et al. (2004); Jin & Qiao (2024); Guo et al. (2024)

### v2

@@ Abstract
We present an analysis of gravitational wave polarization modes within Gravitational Quantum Field Theory (GQFT), a unified theoretical framework reconciling general relativity and quantum field theory. In the six fundamental polarization: two tensors ($\times$), two vectors (x, y), and two scalars (breathing, longitudial) modes of gravitational wave, two tensors and breathing are favored by GQFT. Their distinctive detection signatures in space-based interferometers like LISA and Taiji can be verified. Using first-order orbital dynamics in the Solar System Barycenter frame, we identify three novel observational features: (1) characteristic interference patterns between polarization modes, (2) provide a model-independent response function, and (3) sky-position-dependent optimal detection methods. Our approach provides complete sky coverage through polarization mapping while remaining fully compatible with existing mission designs, notably avoiding the need for challenging direct breathing-mode measurements. The results are presented through comprehensive sky maps, offering both theoretical insights into gravitational wave polarization and practical tools for future detectors. This work establishes a new paradigm for testing fundamental gravity theories through their unique polarization fingerprints.
I Introduction
Gravitational waves (GWs) have been extensively observed through various experiments, with polarization playing a crucial role in exploring extended theories of gravity. In the framework of metric-compatible theories, there are up to six possible polarization modes [1]: two tensor modes (+ and $\times$), two vector modes (x and y), and two scalar modes (breathing and longitudinal). Different gravitational theories predict distinct polarization states. For instance, in general relativity, only the plus and cross modes exist, which have been detected by the LIGO, Virgo, and KAGRA collaborations [2]. In Brans-Dicke gravity [3, 4], an additional breathing mode arises due to the presence of a scalar field. Among various alternative theories, this paper focuses on Gravitational Quantum Field Theory (GQFT) [5, 6, 7, 8].
GQFT was established to reconcile general relativity (GR) and quantum field theory (QFT). It is based on the fundamental principle that the laws of nature are determined by the intrinsic properties of matter’s basic constituents. This principle distinguishes between two types of symmetries: intrinsic symmetries, which are defined by the quantum numbers of quantum fields as elementary particles, and external symmetries, which describe their motion in flat Minkowski spacetime[5, 6]. According to this framework, the intrinsic spin symmetry SP(1,3) must be localized into a spin gauge symmetry following the gauge symmetry principle.
@@ I Introduction
[…]
@@
Beyond its foundational implications, GQFT provides a theoretical framework for exploring a hyperunified field theory capable of unifying all fundamental interactions [10, 11, 12, 13, 14, 15]. Recently, a general theory of the standard model [16] has been developed within the GQFT framework. This theory integrates the two standard models of particle physics and cosmology, offering new insights into the mysteries of the universe’s dark sector. Specifically, it provides fresh perspectives on understanding the nature of dark matter, the dynamics of the early inflationary universe, and the role of dark energy in the current cosmic expansion.
@@
Numerous studies have been conducted within the framework of GQFT, addressing topics such as inflation [17], dark matter [18], and particle physics [19]. The Taiji mission, a space-based gravitational wave (GW) observatory, has been analyzed within the parametrized post-Einsteinian framework to evaluate its capability in detecting different polarization modes. Studies indicate that Taiji can measure both dipole and quadrupole GW emissions [20, 21, 22]. Similarly, the LISA-Tianqin network has been investigated for its sensitivity to the polarization modes of the stochastic GW background [23]. In a general four-dimensional metric theory of gravity, the detection of additional polarization modes beyond the two predicted by GR would suggest the need for an extended theory of gravitation. The specific polarizations observed could help rule out certain theoretical models [24]. Both Taiji [25] and LISA [26] employ a triangular three-detector configuration with equal arm lengths. This paper focuses on the polarization modes in GQFT and their corresponding response characteristics in such a three-detector setup.
@@
The interaction of gravitational waves (GWs) with a detector arm is manifested as fluctuations in the arm length between two detectors in a triangular three-detector configuration. The intensity of the detector’s response depends on the angle between the GW propagation direction and the detector arm. This response is minimal when the GW propagation direction is parallel to the arm and reaches its maximum when the GW direction is perpendicular to the arm. Since gravitational wave sources are distributed across the sky, the angle between the GW propagation direction and the detector arm is not constant. This variation arises due to the orbital motion of space-based detectors, such as Taiji [25] and LISA [26], around the Sun. As these detectors continuously change their positions and orientations, the angle between the GW propagation direction and the detector arm fluctuates, leading to variations in the detector’s sensitivity to GW signals.
@@
Laser interferometers detect gravitational waves (GWs) by measuring the differential length changes between two detector arms. However, maintaining perfectly equal arm lengths over the entire orbital period of a space-based detector is impractical. This challenge necessitates the use of Time Delay Interferometry (TDI) to account for varying arm lengths. Recent advancements in atomic clock technology offer an alternative approach. By precisely measuring the optical path time difference between detectors, these clocks can infer arm-length variations without relying solely on TDI. This method has sparked significant interest in the development of high-precision optical clocks for space-based GW detection[27, 28, 29, 30].
@@
In this paper, we concentrate on investigating novel polarization modes predicted by GQFT that extend beyond GR. Our analysis specifically examines the response of individual detector arms to gravitational wave signals, and will not include considerations of dual-arm interferometric measurements to maintain analytical focus. The paper is organized as follows: it first introduces in §10 the polarization modes in metric theories of gravity, then derives in §II.2 the specific polarization modes within the GQFT framework. The analysis further presents in §II.3 the detector response characteristics in the Solar System Barycenter frame, along with an universal response function under Taiji configuration without specific theory, and providing the optimal observational timing windows for GQFT in §III.
II Methods
II.1 Polarization in metric theory
@@
The action of GW on test masses can be understood through their tidal effects, which manifest as deviations from geodesic motion. This phenomenon is captured by the Jacobi field equation for geodesic deviation, $\dfrac{\mathop{}\!\mathrm{D}^{2}X^{\alpha}}{\mathop{}\!\mathrm{d}\tau^{2}}=T^{\gamma}\nabla_{\gamma}(T^{\delta}\nabla_{\delta}X^{\alpha})=-R^{\alpha}_{\beta\delta\gamma}X^{\delta}T^{\gamma}T^{\beta}.$ In metric theories of gravity, the polarization state is entirely determined by the electric components of the Riemann tensor [31]. When examined in the gravitationally independent flat frame (the natural framework for GQFT), this description simplifies to the form, $\dfrac{\mathop{}\!\mathrm{d}^{2}X^{i}}{\mathop{}\!\mathrm{d}\tau^{2}}=-R^{i}_{0j0}X^{j}$, which describes weak, plane, null gravitational waves [1]. The symmetry properties of the Riemann tensor reveal that $R^{i}_{0j0}$ possesses six independent components, indicating that gravitational waves can exhibit up to six distinct polarization modes.
@@
For a wave propagating along the z-direction, we define $R_{i0j0}(t)$ as the “driving force matrix” $S_{ij}(t)$, a 3-dimensional symmetric matrix. The polarization modes can be further classified using the Newman-Penrose formalism [32]. Working with tetrad indices $(l,n,m,\bar{m})$, we consider tensor contractions with the corresponding null vectors, $X_{efgh}:=X_{\mu\nu\rho\sigma}e^{\mu}f^{\nu}g^{\rho}h^{\sigma}$, where $(e,f,g,h)$ range over the null tetrad basis $(l,n,m,\bar{m})$. Taking a null tetrad frame as follows:
$l=\frac{1}{\sqrt{2}}(\partial_{t}+\partial_{z}),\quad n=\frac{1}{\sqrt{2}}(\partial_{t}-\partial_{z})\\ m=\frac{1}{\sqrt{2}}(\partial_{x}+i\partial_{y}),\quad\bar{m}=\frac{1}{\sqrt{2}}(\partial_{x}-i\partial_{y}).$ (1)
@@
In the null frame, where the field depends solely on retarded time (tz)(t-z), the Riemann tensor satisfies the condition, $R_{efgh,p}=0$. Here, $(e,f,g,h)$ span the null tetrad $(l,n,m,\bar{m})$, while $(p,q)$ range over $(l,m,\bar{m})$. Substituting these into the differential Bianchi identity yields, $R_{ef[pq,n]}={\dfrac{1}{3}R_{efpq,n}=0}$. This implies that, up to trivial constants, we must have: $R_{efpq}=R_{pqef}=0$. Consequently, the only non-vanishing components of the Riemann tensor are of the form $R_{pnqn}$.
@@
Under the null condition, the original 12 NP scalars reduce to just 4 independent quantities: $\Psi_{2},\Psi_{3},\Psi_{4},\Phi_{22}$. Two real scalars $\Psi_{2},\Phi_{22}$, and Two complex scalars $\Psi_{3},\Psi_{4}$. These remaining NP null scalars collectively contain 6 independent degrees of freedom, corresponding exactly to the degrees of freedom in the Riemann tensor for weak, plane, null gravitational waves [33]:
@@
where $A$ ranges over $(+,\times,x,y,b,l),$ covering 6 polarization modes.
@@ II.2 Polarization in GQFT
Dynamical analysis of GQFT identifies five fundamental degrees of freedom in the gravitational sector [9]. Although this implies the theoretical existence of five gravitational wave polarization modes, only three physical polarizations emerge as observationally significant when analyzing their tidal effects on test masses. The geodesic deviation equation, which governs observable gravitational wave effects, exclusively captures tidal forces. Consequently, the remaining two vector-mode polarizations remain observationally inaccessible within this framework, as they encode intrinsic spin properties and represent manifestations of equivalence principle violations rather than measurable tidal deformations.
@@
$\begin{bmatrix}~&~&~&~\\ ~&\hat{h}_{+}&\hat{h}_{\times}&~\\ ~&\hat{h}_{\times}&-\hat{h}_{+}&~\\ ~&~&~&~\end{bmatrix}$, (6)
@@
II.3 Response on the detectors
As mentioned above, GW propagation can be expressed as a linear combination of different polarization tensors [34],
@@ II.3 Response on the detectors
Nearly monochromatic gravitational waves are generated under the quadrupole approximation[35]. For HM Cancri, refer to the waveform presented [36, 37]
$h_{+}(t)$ $=\frac{2(G\mathcal{M})^{5/3}(\pi f)^{2/3}}{d_{L}c^{4}}(1+\cos^{2}\iota)\cos\Phi(t),$ (14)
$h_{\times}(t)$ $=\frac{4(G\mathcal{M})^{5/3}(\pi f)^{2/3}}{d_{L}c^{4}}\cos\iota\sin\Phi(t).$ (15)
The processing of b mode shall follow the procedures of tensor mode. A comparative analysis of GQFT wave equations reveals that the b-mode and tensor modes differ by a characteristic factor $(1+\gamma_{W})$, where $\gamma_{W}$ represents a fundamental GQFT parameter characterized by the mass ratio $\gamma_{W}=M_{\mathcal{A}}^{2}/\bar{M}_{\kappa}^{2}$ [5, 6, 7, 9, 8, 16]. Here, $M_{\mathcal{A}}$ is the mass of spin gauge boson and $\bar{M}_{\kappa}$ denotes the fundamental mass at Planck scale. Indeed, the polarization of b mode is generally considered to be very small compared to pure geometry Theory, but in GQFT it has a contribution from the spin gauge field, so it can reach the order of tensor mode [38].
Simultaneously setting the orbital inclination and initial phase to zero also better unifies the consistency of waveform amplitude. We employed the following waveform here
$h_{+}(t)$ $=\frac{4(G\mathcal{M})^{5/3}(\pi f)^{2/3}}{d_{L}c^{4}}\cos\Phi(t),$ (16)
$h_{\times}(t)$ $=\frac{4(G\mathcal{M})^{5/3}(\pi f)^{2/3}}{d_{L}c^{4}}\sin\Phi(t),$ (17)
$h_{b}(t)$ $=\frac{2(1+\gamma_{W})(G\mathcal{M})^{5/3}(\pi f)^{2/3}}{d_{L}c^{4}}\cos\Phi(t).$ (18)
Here, $d_{L}$ is the luminosity distance, $\mathcal{M}=M\eta^{\frac{3}{5}}$ is the chirp mass with $\eta=\dfrac{m_{1}m_{2}}{M^{2}}$ and $M=m_{1}+m_{2}$ is the total mass. Meanwhile, with regard to phase, we consider a near-monochromatic chirp signal, which means $\Phi(t)=2\pi ft+\pi\dot{f}t^{2}+\phi_{0}$, where $\dot{f}=\dfrac{96}{5}\pi^{8/3}\left(\dfrac{G\mathcal{M}}{c^{3}}\right)^{5/3}f^{11/3}.$ Since this work aims to develop a polarization-focused model for distinguishing between GQFT and GR, the analysis remains independent of specific waveform models, thereby maintaining generality.
[…]
@@
When examining the effect of GWs on the detector, consider a photon emitted from spacecraft $i$ at time $t_{i}$ and received at spacecraft $j$ at time $t_{j}$. The nominal arm length is $\ell_{ij}$, and its variation under low-frequency GWs (where the transfer function approaches unity) can be expressed as [39, 36, 40]